### v1

@@ Abstract
The anisotropy of the redshift space bispectrum depends upon the orientation of the triangles formed by three ${\bm{k}}$ modes with respect to the line of sight. For a triangle of fixed size ($k_{1}$) and shape ($\mu,t$), this orientation dependence can be quantified in terms of angular multipoles $B_{\ell}^{m}(k_{1},\mu,t)$ which contain a wealth of cosmological information. We propose a fast and efficient FFT-based estimator that computes bispectrum multipole moments $B_{\ell}^{m}$ of a 3D cosmological field for all possible $\ell$ and $m$ (including $m\neq 0$). The time required by the estimator to compute all multipoles from a gridded data cube of volume $N_{g}^{3}$ scales as $\sim N_{g}^{3}\log{(N_{g})}$ ) in contrast to the direct computation technique which requires time $\sim N_{g}^{6}$. Here, we demonstrate the formalism and validate othe estimator using a simulated non-Gaussian field for which the analytical expressions for all bispectrum multipoles are known. The estimated results are found to be in good agreement with the analytical predictions for all $16$ non-zero multipoles (up to $\ell=6,m=6$). We expect the $m\neq 0$ bispectrum multipoles to significantly enhance the information available from galaxy redshift surveys, and future redshifted 21-cm observations.
@@ I Introduction
The spatial distribution of galaxies has emerged as one of the primary probes of large-scale structure (LSS) of the Universe. The simplest model of inflation predicts the primordial density fluctuations that seed the LSS to be a Gaussian random field. Nevertheless, the subsequent non-linear growth and non-linear biasing induce non-Gaussianity in the galaxy distribution. Further, many inflationary paradigms speculate non-Gaussian primordial fluctuations. In any scenario, it is essential to exploit the statistics that capture the non-Gaussian nature of galaxy distribution and utilize the maximum potential of data from surveys such as SDSS [^1] https://www.sdss.org/ [1], DESI [^2] https://www.desi.lbl.gov/ [2], LSST [^3] https://www.lsst.org/ [3], and EUCLID [^4] https://www.euclid-ec.org/ [4]. Bispectrum is the lowest-order statistic sensitive to non-Gaussianity in the underlying distribution. It is a function of closed triangles formed by three ${\bm{k}}$ vectors and contains the information on mode coupling. The inclusion of bispectrum significantly supplements our comprehension of the standard model of cosmology, enabling robust constraints on the primordial non-Gaussianity [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], modeling the inflation [16, 17, 18], lifting the degeneracy between $\Omega_{m}$ and $b_{1}$ [19], constraining cosmological parameters [20, 21, 22], and galaxy bias parameters [23].
@@
The observed galaxy bispectrum is anisotropic along the line of sight direction caused by the peculiar motion of galaxies. This anisotropy of the redshift space bispectrum can be studied by decomposing it into spherical harmonic $Y_{\ell}^{m}(\hat{p})$ basis [24, 25, 26]. The expansion coefficients in $Y_{\ell}^{m}(\hat{p})$ basis, bispectrum multipole moments $B_{\ell}^{m}$, can be measured from the survey data. While the majority of work has focused on monopole moment $B_{0}^{0}$ only [21], the measurements of few higher order multipoles have also been utilized to constrain cosmologies [27, 28, 29, 20, 22]. To the best of our knowledge, all of the previous studies have been restricted to $m=0$ multipoles. [20] have shown that including $\ell=2$ and 4 with $m=0$ in the analysis have only marginal effect on the cosmological parameter posteriors. A recent study by [30] has shown that the incorporation of $m\neq 0$ moments significantly improves the information gain. It is crucial to accurately measure the higher-order multipoles with $m\neq 0$ for precision cosmology.
Following the theoretical framework introduced by [25] and [26], in this paper, we present a fast and accurate estimator to calculate the bispectrum multipoles $B_{\ell}^{m}$, including those with $m\neq 0$, considering triangles of all possible unique shapes.
@@ II Methodology
and orientation with respect to the LoS direction $\hat{z}$ by a unit vector $\hat{\bm{p}}$. The cosine of the angle between the three sides of the triangle and $\hat{z}$ are respectively given by
@@
where the integration is over $4\pi$ steradian, corresponding to all possible orientations of a triangle of a fixed size $(k_{1})$ and shape $(\mu,t)$. From the observational point of view, we consider a finite volume $V$ with $N_{g}^{3}$ grid points. The multipole moments can be calculated by identifying the set $\mathcal{T}$ of all triangles of a fixed size $(k_{1})$ and shape $(\mu,t)$, and using
$\bar{B}_{\ell}^{m}(k_{1},\mu,t)=\sqrt{\frac{2\ell+1}{4\pi}}\frac{\sum_{\mathcal{T}}[Y_{\ell}^{m}(\hat{\bm{p}})]^{*}B^{s}(\hat{\bm{p}},k_{1},\mu,t)}{\sum_{\mathcal{T}}|Y_{\ell}^{m}(\hat{\bm{p}})|^{2}}~{},$ (5)
@@
where $\sum_{\mathcal{T}}$ denotes a sum over all triangle orientations that are accessible in the 3D gridded data cube. The multipole moments $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ (Eq.4) do not exactly match $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ (Eq.5) because of the discrete sampling of ${\bm{k}}$ space whereby the integration is replaced with a summation [20, 31]. We have accounted for this effect and will describe it in the next section.
@@
Figure 1: The binning scheme of the estimator. This demonstrates the division of ${\bm{k}}$ space into multiple planes perpendicular to the Line of Sight (LoS) direction $\hat{z}$, with each plane further subdivided into concentric annular rings. We show a closed triangle formed by three modes ${\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3}$ that fall in different annular rings labeled as $a_{1},a_{2},a_{3}$. The combination of three rings and planes defines a bin of triangles that have a fixed shape ($\mu,t$), size ($k_{1}$), and orientation ($\hat{\bm{p}}$) with respect to LoS. If this triangle is rotated around the LoS such that three ${\bm{k}}$ modes are restricted within respective annular rings, then the rotated triangle will still have the same orientation ($\hat{\bm{p}}$) relative to LoS as that of the original triangle. The bin contains all such triangles. The ${\bm{k}}$ vectors are decomposed into their $k_{\parallel}\hat{z}$ and ${\bm{k}}_{\perp}$ components with respect to LoS. Note that the middle plane is considered to be centered at the origin ${\bm{k}}=0$ (hence $k_{2\parallel}=0$) to avoid complexity in the illustration.
@@
The direct implication of Eq.5 is immensely time-consuming as it involves identifying triangles, which requires computations of order $N_{g}^{6}$. Several earlier works by [32, 33, 13] have proposed an FFT based technique to compute binned statistics in Fourier space, which considerably speeds up the bispectrum computation. Following a similar formalism, [34] has proposed a fast estimator that considers spherical shells in ${\bm{k}}$ space to compute the binned bispectrum monopole $\bar{B}_{0}^{0}(k_{1},\mu,t)$. Further, this can be easily extended to estimate the higher multipoles with $m=0$, as has been implemented in [35]. However, the $m\neq 0$ multipoles introduce additional complexity due to $e^{im\phi}$ phase factor in the spherical harmonics, which encapsulates the dependence on the azimuthal direction. It is not possible to incorporate this phase factor if we use spherical shells in ${\bm{k}}$ space to compute the binned bispectrum. We have overcome this difficulty by considering a different binning scheme, as shown in Fig.1. Here the box can be viewed as a collection of $N_{g}$ 2D planes normal to $\hat{z}$. Each plane is uniformly divided into $N_{r}$ concentric annular rings of width $\delta k_{\perp}$, and three such planes are shown in Fig.1. Decomposing a ${\bm{k}}$ vector into components ${\bm{k}}_{\perp}$ and $k_{\parallel}$ that are respectively perpendicular and parallel to $\hat{z}$, the closed triangle condition yields to two conditions $k_{1\parallel}+k_{2\parallel}+k_{3\parallel}=0$ and ${\bm{k}}_{1\perp}+{\bm{k}}_{2\perp}+{\bm{k}}_{3\perp}=0$. The first condition constrains the choice of the three planes corresponding to $(k_{1\parallel},k_{2\parallel},k_{3\parallel})$. Here, we independently fix two planes namely $k_{1\parallel}$ and $k_{2\parallel}$, for which the third plane is determined using $k_{3\parallel}=-k_{1\parallel}-k_{2\parallel}$.
@@
According to the second condition, the perpendicular components of the three wave vectors must form a closed triangle on the 2D plane perpendicular to $\hat{z}$. To estimate the binned bispectrum, we consider three annular rings labeled $(a_{1},a_{2},a_{3})$ respectively located in the planes $(k_{1\parallel},k_{2\parallel},k_{3\parallel})$. We define the binned bispectrum estimator as
$\hat{B}^{s}$ $(k_{1\perp},k_{2\perp},k_{3\perp},k_{1\parallel},k_{2\parallel})$ (6)
$=\frac{1}{V~{}N_{\rm tri}}\sum_{{\bm{k}}_{1\perp}}\sum_{{\bm{k}}_{2\perp}}\sum_{{\bm{k}}_{3\perp}}\Big{[}\Delta({\bm{k}}_{1\perp},k_{1\parallel})~{}\Delta({\bm{k}}_{2\perp},k_{2\parallel})$
×$\quad\times\Delta({\bm{k}}_{3\perp},k_{3\parallel})\Big{]}~{}\delta_{\rm K}({\bm{k}}_{1\perp}+{\bm{k}}_{2\perp}+{\bm{k}}_{3\perp})\,,$
@@
where $({\bm{k}}_{1\perp},{\bm{k}}_{2\perp},{\bm{k}}_{3\perp})$ refer to the possible ${\bm{k}}_{\perp}$ modes in $(a_{1},a_{2},a_{3})$ respectively. The 2D Kronecker delta function $\delta_{\rm K}({\bm{k}}_{1\perp}+{\bm{k}}_{2\perp}+{\bm{k}}_{3\perp})$ ensures that the estimator picks up a contribution only when the three vectors $({\bm{k}}_{1\perp},{\bm{k}}_{2\perp},{\bm{k}}_{3\perp})$ form a closed triangle, and $N_{\rm tri}$ is the total number of such triangles. Here, $(k_{1\perp},k_{2\perp},k_{3\perp})$ respectively refer to the mean wave number of the ${\bm{k}}_{\perp}$ modes in $(a_{1},a_{2},a_{3})$ respectively.
@@
$\delta_{\rm K}({\bm{k}}_{1\perp}+{\bm{k}}_{2\perp}+{\bm{k}}_{3\perp})=\frac{1}{N_{\rm g}^{2}}\sum_{\bm{x}_{\perp}}\exp(-i[{\bm{k}}_{1\perp}+{\bm{k}}_{2\perp}+{\bm{k}}_{3\perp}]\cdot\bm{x}_{\perp})~{}.$ (7)
@@
$\hat{B}^{s}$ $(k_{1\perp},k_{2\perp},k_{3\perp},k_{1\parallel},k_{2\parallel})$ (8)
$=\frac{1}{V~{}N_{\rm tri}}\frac{1}{N_{\rm g}^{2}}\sum_{\bm{x}_{\perp}}D(k_{1\perp},k_{1\parallel},\bm{x}_{\perp})D(k_{2\perp},k_{2\parallel},\bm{x}_{\perp})$
×$\quad\times D(k_{3\perp},k_{3\parallel},\bm{x}_{\perp})~{},$
@@
$D(k_{1\perp},k_{1\parallel},\bm{x}_{\perp})=\sum_{{\bm{k}}_{1\perp}}\Delta({\bm{k}}_{1\perp},k_{1\parallel})\exp(-i{\bm{k}}_{1\perp}\cdot\bm{x}_{\perp})~{},$ (9)
which is the inverse FFT of the field $\Delta({\bm{k}}_{1\perp},k_{1\parallel})$ restricted to the annular ring $a_{1}$. The other $D$s are similarly defined considering $a_{2}$ and $a_{3}$.
@@
$N_{\rm tri}=\sum_{{\bm{k}}_{1\perp}}\sum_{{\bm{k}}_{2\perp}}\sum_{{\bm{k}}_{3\perp}}\delta_{\rm K}({\bm{k}}_{1\perp}+{\bm{k}}_{2\perp}+{\bm{k}}_{3\perp})~{}.$ (10)
@@
$N_{\rm tri}=\frac{1}{N_{\rm g}^{2}}\sum_{\bm{x}_{\perp}}I(k_{1\perp},\bm{x}_{\perp})I(k_{2\perp},\bm{x}_{\perp})I(k_{3\perp},\bm{x}_{\perp})~{},$ (11)
@@
$I(k_{1\perp},\bm{x}_{\perp})=\sum_{{\bm{k}}_{1\perp}}\exp(-i{\bm{k}}_{1\perp}\cdot\bm{x}_{\perp})\,,$ (12)
and the other $I$s are similarly defined using $a_{2}$ and $a_{3}$.

### v2

@@ Abstract
The anisotropy of the redshift space bispectrum depends upon the orientation of the triangles formed by three ${\bm{k}}$ modes with respect to the line of sight. For a triangle of fixed size ($k_{1}$) and shape ($\mu,t$), this orientation dependence can be quantified in terms of angular multipoles $B_{\ell}^{m}(k_{1},\mu,t)$ which contain a wealth of cosmological information. We propose a fast and efficient FFT-based estimator that computes the bispectrum multipole moments $B_{\ell}^{m}$ of a 3D cosmological field for all possible $\ell$ and $m$ (including $m\neq 0$). The time required by the estimator to compute all multipoles from a gridded data cube of volume $N_{g}^{3}$ scales as $\sim\mathcal{O}(N_{g}^{4})$ in contrast to the direct computation technique which requires time $\sim\mathcal{O}(N_{g}^{6})$. Here, we demonstrate the formalism and validate the estimator using a simulated non-Gaussian field for which the analytical expressions for all the bispectrum multipoles are known. The estimated results are found to be in good agreement with the analytical predictions for all $16$ non-zero multipoles (up to $\ell=6,m=6$). We expect the $m\neq 0$ bispectrum multipoles to significantly enhance the information available from galaxy redshift surveys and future redshifted 21-cm observations.
@@ I Introduction
The spatial distribution of galaxies has emerged as one of the primary probes of large-scale structure (LSS) of the Universe. The simplest model of inflation predicts the primordial density fluctuations that seed the LSS to be a Gaussian random field. Nevertheless, the subsequent non-linear growth and non-linear biasing induce non-Gaussianity in the galaxy distribution. Further, many inflationary paradigms speculate non-Gaussian primordial fluctuations. In any scenario, it is essential to exploit the statistics that capture the non-Gaussian nature of galaxy distribution and utilize the maximum potential of data from surveys such as SDSS [^1] https://www.sdss.org/ [1], DESI [^2] https://www.desi.lbl.gov/ [2], LSST [^3] https://www.lsst.org/ [3], and EUCLID [^4] https://www.euclid-ec.org/ [4]. The bispectrum is the lowest-order statistic sensitive to non-Gaussianity in the underlying distribution. It is a function of the closed triangles formed by three ${\bm{k}}$ vectors and contains the information on mode coupling. The inclusion of the bispectrum significantly supplements our comprehension of the standard model of cosmology, enabling robust constraints on the primordial non-Gaussianity [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], modeling the inflation [16, 17, 18], lifting the degeneracy between $\Omega_{m}$ and $b_{1}$ [19], constraining cosmological parameters [20, 21, 22], and galaxy bias parameters [23].
@@
The observed galaxy bispectrum is anisotropic along the line of sight (LoS) direction because of the peculiar motion of galaxies. This anisotropy of the redshift space bispectrum can be studied by decomposing it into spherical harmonics $Y_{\ell}^{m}(\hat{p})$ [24, 25, 26]. The expansion coefficients in $Y_{\ell}^{m}(\hat{p})$ basis, the bispectrum multipole moments $B_{\ell}^{m}$, can be measured from the survey data. While the majority of work has focused on the monopole moment $B_{0}^{0}$ only (e.g. [21]), a few have considered the higher order multipoles [27, 28, 20, 22]. Ivanov et al. [20] have shown that including the $\ell=2$ and $4$ multipoles in the analysis only leads to a marginal improvement of the cosmological parameters. However, the studies mentioned above have all been restricted to the $m=0$ multipoles. Gagrani and Samushia [29] has studied the information loss associated with restricting the study to $m=0$ multipoles using the Fisher information formalism and find that $m\neq 0$ multipoles do not carry significant information. However, in a recent work, Heinrich et al. [30] has shown that incorporating the $m\neq 0$ multipole moments significantly improves the information gain. It is crucial to accurately measure the higher-order multipoles with $m\neq 0$ for precision cosmology. Nan et al. [31] have used the halo model to make analytical predictions for all the non-zero higher-order multipoles (including those with $m\neq 0$) up to $\ell=4$, $m=4$, and compared these with the values computed from mock galaxy samples. In this paper, we present a fast and accurate estimator to calculate the bispectrum multipoles $B_{\ell}^{m}$, including those with $m\neq 0$, considering triangles of all possible unique shapes. To the best of our knowledge, this is the first implementation of an FFT-based fast bispectrum estimator that includes the $m\neq 0$ multipoles. This paper builds on the formalism and bispectrum parametrization presented in [25] and [26].
@@ II Methodology
Here we opt for the parametrization prescribed by [25]. For a triangle with $k_{1}\geq k_{2}\geq k_{3}$, its size is parameterized by $k_{1}=|{\bm{k}}_{1}|$, shape by two dimensionless parameters
@@
and orientation with respect to $\hat{z}$ by a unit vector $\hat{\bm{p}}$. The cosine of the angle between the three sides of the triangle and $\hat{z}$ are respectively given by
@@
where the integration is over $4\pi$ steradian, corresponding to all possible orientations of a triangle of a fixed size $(k_{1})$ and shape $(\mu,t)$. The integration assumes the continuum limit for the ${\bm{k}}$ modes. Hereafter, we refer to the results obtained by analytically performing this integral as $[\bar{B}_{\ell}^{m}]_{\rm c}$. From the observational point of view, we consider a finite volume $V$ with $N_{g}^{3}$ grid points. The multipole moments can be calculated by identifying the set $\mathcal{T}$ of all triangles of a fixed size $(k_{1})$ and shape $(\mu,t)$, and using
$[\bar{B}_{\ell}^{m}]_{\rm d}(k_{1},\mu,t)=\sqrt{\frac{2\ell+1}{4\pi}}\frac{\sum_{\mathcal{T}}[Y_{\ell}^{m}(\hat{\bm{p}})]^{*}B^{s}(\hat{\bm{p}},k_{1},\mu,t)}{\sum_{\mathcal{T}}|Y_{\ell}^{m}(\hat{\bm{p}})|^{2}}~{},$ (5)
@@
where $\sum_{\mathcal{T}}$ denotes a sum over all triangle orientations that are accessible in the 3D gridded data cube. The multipole moments $[\bar{B}_{\ell}^{m}]_{\rm c}$ (Eq. 4) do not exactly match $[\bar{B}_{\ell}^{m}]_{\rm d}$ (Eq. 5) because of the discrete sampling of ${\bm{k}}$ space whereby the integration is replaced with a summation [20, 32]. We have accounted for this effect and will describe it in the next section.
@@
Figure 1: The binning scheme of the estimator. This demonstrates the division of ${\bm{k}}$ space into multiple planes perpendicular to the Line of Sight (LoS) direction $\hat{z}$, with each plane further subdivided into concentric annular rings. We show a closed triangle formed by three modes ${\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3}$ that fall in different annular rings labeled as $r_{1},r_{2},r_{3}$. The combination of three rings ($r_{1},r_{2},r_{3}$) and planes ($n_{1},n_{2},n_{3}$) defines a bin of triangles that have a fixed shape ($\mu,t$), size ($k_{1}$), and orientation ($\hat{\bm{p}}$) with respect to the LoS. If this triangle is rotated around the LoS such that three ${\bm{k}}$ modes are restricted within respective annular rings, then the rotated triangle will still have the same orientation ($\hat{\bm{p}}$) relative to the LoS as that of the original triangle. The bin contains all such triangles. The ${\bm{k}}$ vectors are decomposed into their $k_{\parallel}\hat{z}$ and ${\bm{k}}_{\perp}$ components with respect to the LoS. Note that the middle plane is considered to be centered at the origin ${\bm{k}}=0$ (hence $k_{\parallel n_{2}}=0$) to avoid complexity in the illustration.
@@
The direct implementation of Eq. 5 is immensely time-consuming as it involves identifying triangles, which requires computations of order $N_{g}^{6}$. Several earlier works by [33, 34, 13] have proposed an FFT based technique to compute binned statistics in Fourier space, which considerably speeds up the bispectrum computation. Following a similar formalism, [35] has proposed a fast estimator that considers spherical shells in ${\bm{k}}$ space to compute the binned bispectrum monopole $\bar{B}_{0}^{0}(k_{1},\mu,t)$. Further, this can be easily extended to estimate the higher multipoles with $m=0$, as has been implemented in [36]. However, the $m\neq 0$ multipoles introduce additional complexity due to $e^{im\phi}$ phase factor in the spherical harmonics, which encapsulates the dependence on the azimuthal direction. It is not possible to incorporate this phase factor if we use spherical shells in ${\bm{k}}$ space to compute the binned bispectrum. We have overcome this difficulty by considering a different binning scheme, as shown in Fig. 1. Here, the box can be viewed as a collection of $N_{g}$ 2D planes normal to $\hat{z}$, which we label using $n=1,2,3,...,N_{g}$ each with a corresponding $k_{\parallel n}$. Each plane is divided into $N_{r}$ concentric annular rings, which we label using $r=1,2,3,...,N_{r}$. Considering the annular ring $r$, $\{r\}$ denotes the set of ${\bm{k}}_{\perp}$ modes contained in the ring, $k_{\perp r}$ denotes its mean radius and $(\delta k_{\perp})_{r}$ its width. The largest ring is restricted to a maximum outer radius of $N_{g}/3$ by the periodic boundary condition imposed on the finite volume [35]. The annular rings and three planes are schematically shown in Fig. 1.
@@
To estimate the binned bispectrum, we consider three annular rings $(r_{1},r_{2},r_{3})$ respectively located in the planes $(n_{1},n_{2},n_{3})$. As discussed earlier, only $n_{1}$ and $n_{2}$ are independently specified, $n_{3}$ is determined from the condition $k_{\parallel n_{1}}+k_{\parallel n_{2}}+k_{\parallel n_{3}}=0$. We define the binned bispectrum estimator as
$\hat{B}^{s}(k_{\perp r_{1}},k_{\perp r_{2}},k_{\perp r_{3}},k_{\parallel n_{1}},k_{\parallel n_{2}})=\frac{1}{V~{}N_{\rm tri}}\sum_{{\bm{u}}}\sum_{{\bm{v}}}\times$ (6)
$\sum_{{\bm{w}}}\Delta({\bm{u}},k_{\parallel n_{1}})\Delta({\bm{v}},k_{\parallel n_{2}})\Delta({\bm{w}},k_{\parallel n_{3}})\,\delta_{\rm K}({\bm{u}}+{\bm{v}}+{\bm{w}}),$
@@
where ${\bm{u}}\in\{r_{1}\},{\bm{v}}\in\{r_{2}\}$ and ${\bm{w}}\in\{r_{3}\}$ refer to three 2D wave vectors in the plane perpendicular to $\hat{z}$. The 2D Kronecker delta function $\delta_{\rm K}({\bm{u}}+{\bm{v}}+{\bm{w}})$ ensures that the estimator receives a contribution only when the three vectors $({\bm{u}},{\bm{v}},{\bm{w}})$ form a closed triangle and $N_{\rm tri}$ is the total number of such triangles. The binned bispectrum estimator (eq. 6) provides an estimate of the bispectrum averaged over a set of triangles that have nearly the same shape, size, and orientation with respect to $\hat{z}$.
@@
$\delta_{\rm K}({\bm{u}}+{\bm{v}}+{\bm{w}})$ $=\frac{1}{N_{\rm g}^{2}}\sum_{\bm{x}_{\perp}}\exp(-i[{\bm{u}}+{\bm{v}}+{\bm{w}}]\cdot\bm{x}_{\perp})~{}.$ (7)
@@
$\hat{B}^{s}(k_{\perp r_{1}},k_{\perp r_{2}},k_{\perp r_{3}},k_{\parallel n_{1}},k_{\parallel n_{2}})=\frac{1}{V~{}N_{\rm tri}}\frac{1}{N_{\rm g}^{2}}\times$ (8)
$\quad\sum_{\bm{x}_{\perp}}D(r_{1},n_{1},\bm{x}_{\perp})D(r_{2},n_{2},\bm{x}_{\perp})D(r_{3},n_{3},\bm{x}_{\perp})~{},$
@@
$D(r_{1},n_{1},\bm{x}_{\perp})=\sum_{{\bm{u}}\in\{r_{1}\}}\Delta({\bm{u}},k_{\parallel n_{1}})\exp(-i{\bm{u}}\cdot\bm{x}_{\perp})~{},$ (9)
which is the 2D inverse FFT of the field $\Delta({\bm{u}},k_{\parallel n_{1}})$ with ${\bm{u}}$ restricted to the annular ring $r_{1}$. The other $D$s are similarly defined.
@@
$N_{\rm tri}=\sum_{{\bm{u}}\in\{r_{1}\}}\sum_{{\bm{v}}\in\{r_{2}\}}\sum_{{\bm{w}}\in\{r_{3}\}}\delta_{\rm K}({\bm{u}}+{\bm{v}}+{\bm{w}})~{}.$ (10)
@@
$N_{\rm tri}=\frac{1}{N_{\rm g}^{2}}\sum_{\bm{x}_{\perp}}I(r_{1},\bm{x}_{\perp})I(r_{2},\bm{x}_{\perp})I(r_{3},\bm{x}_{\perp})~{},$ (11)
@@
$I(r_{1},\bm{x}_{\perp})=\sum_{{\bm{u}}\in\{r_{1}\}}\exp(-i{\bm{u}}\cdot\bm{x}_{\perp})~{},$ (12)
and the other $I$s are similarly defined.